# Leveraging the Bias-Variance Tradeoff in Quantum Chemistry for Accurate Negative Singlet-Triplet Gap Predictions: A Case for Double-Hybrid DFT

### v1

@@ 1 Introduction
The energy criterion for organic light-emitting diodes (OLEDs) to operate via thermally activated delayed fluorescence (TADF) is that the energy of the first excited singlet state (S1) must be no more than $0.1$ eV higher than that of the triplet state (T1)[1, 2]. This energy criterion facilitates thermally aided reverse intersystem crossing (RISC), enabling the theoretical complete transfer of the T1 population to the emissive state, S1[3]. A derivative of the triangular-shaped molecule, heptazine (aka heptaazaphenalene, 7AP), with a central nitrogen (N) atom in an anti-aromatic 12-annulene framework, showed delayed fluorescence without thermal assistance, indicating a negative S1–T1 gap (STG)[4]. This mechanism paves the way for designing OLEDs that leverage exothermic delayed fluorescence from inverted singlet and triplet excited states (DFIST). Direct spectroscopic evidence of a negative STG ($-0.047\pm 0.007$ eV) was obtained for pentaazaphenalene (5AP) using anion photoelectron spectroscopy and fluorescence measurements [5]. Negative STG was also was also observed in the transient photoluminescence data of dialkylamine-substituted 5AP[6]. Computational studies have further suggested that azaphenalenes (APs) with other substitution patterns, as well as the boron (B) analog of 1AP, boraphenalene (1BP), also exhibit negative STGs[7, 8, 9, 10, 11, 12, 13, 14]. Non-alternant hydrocarbons and their substituted analogs have been shown to have the potential to exhibit negative STGs[15, 16, 17, 18, 19]. Notably, substituted analogs of the non-fused bicyclic hydrocarbon have demonstrated negative STGs, attributed to through-bond charge-transfer states[20].
@@
A comprehensive search across $12,880$ structurally diverse small molecules revealed no exceptions to Hund’s rule, indicating that achieving the electronic structure criteria for ${\rm STG}<0$ requires nontrivial molecular structures[21]. Moreover, the typical magnitudes of negative STG are much smaller or similar to the average errors associated with popular quantum chemistry and density functional theory approximations (DFAs). Consequently, due to the rarity of molecules with negative STGs and the limitations of current quantum chemical approximations, accurately identifying molecules exhibiting ${\rm STG}<0$ via high-throughput screening—with minimal false-positive and false-negative predictions—remains challenging. The linear-response time-dependent density functional theory (LR-TD-DFT) formalism provides qualitative accuracy—i.e., gets the sign of the STG for DFIST candidates correct—only within the double-hybrid (dh) DFT framework containing many-body correlation via the second-order perturbation theory (MP2). With the right combination of ingredients, dh-DFT can reliably predict a wide range of molecular properties[22, 23, 24, 23]. Nevertheless, selecting suitable exchange-correlation (XC) functionals within the dh-DFT framework remains challenging, especially for novel problems lacking sufficient reference data.
@@
We present a scheme that leverages the bias-variance tradeoff within a single DFT formalism. To achieve this, we evaluate various computational methods for predicting twelve STGs in triangular molecules, benchmarking them against previously reported theoretical best estimates[14]. We examine dh-DFAs exhibiting very low variance for STG predictions, and reparametrize them to minimize the bias. These optimal low-bias models are used as internal references to correct the predictions of the low-variance models. The internally referenced scaling approach is demonstrated by predicting STGs with low bias and low variance, resulting in excellent overall accuracy. We discuss the merits and drawbacks of this approach for its further applications.
@@ 2 Dataset and Computational Methods
We consider twelve systems (see Figure 1) using equilibrium geometries from Ref. [14], determined at the CCSD(T) level with the frozen core approximation and the cc-pVTZ basis set. Using these geometries, we performed single-point excited state calculations with various methods. In addition, we collected TBEs of the S1 and T1 energies from Ref. [14] to serve as reference values for benchmarking. As stated in Ref. [14], structure 11 in Figure 1 (with $D_{3\rm h}$ symmetry) corresponds to a saddle point on the potential energy surface at the MP2/6-311G($d$,$p$) level, whereas the true minimum geometry exhibits $C_{3\rm h}$ symmetry (structure 12). Similarly, a separate study[25] found that the high-symmetry forms of structures 6, 7, and 8 are not true minima and structure 2 is a very shallow potential well at the CCSD(T)/cc-pVTZ level[25]. Overall, four of the twelve structures are transition states; however, we include them in our benchmarking of STGs due to the availability of TBE-level results and the lack of alternative high-quality data.
@@
Laplace-transformed, density-fitted, local versions of correlated methods—CC2 [26] and ADC(2) [27]—offer attractive speedups for modeling excited states of medium-sized molecules. We denote these methods as L-CC2 and L-ADC(2), and performed calculations with the cc-pVDZ, aug-cc-pVDZ, cc-pVTZ, and aug-cc-pVTZ basis sets, using the corresponding JKFIT and MP2FIT auxiliary basis sets. L-CC2 and L-ADC(2) are the same methods denoted as LT-DF-CC2 and LT-DF-ADC(2) elsewhere. L-CC2 and L-ADC(2) calculations were conducted using Molpro (version 2015.1) [28]. L-CC2 has been show to deviate from the canonical variant CC2 by $\leq 0.05$ eV, irrespective of the character of the excited states, including that of charge-transfer states for which CC2 can have huge error [29]. The systems studied in this work do not have charger transfer characteristics, hence CC2 delivers good accuracy for STGs in agreement with TBEs[14]. This implementation has demonstrated excellent accuracy in modeling the excited states of BODIPY derivatives [30, 31]. For comparison, we collected S1 and T1 energies obtained using the linear-response coupled-cluster methods, EOM-CCSD and CC3, with the aug-cc-pVDZ basis set from Ref. [14]. Additionally, we performed ADC(2) calculations using the cc-pVDZ, cc-pVTZ, aug-cc-pVDZ, and aug-cc-pVTZ basis sets with the resolution-of-the-identity (RI) approximation [32, 33] as implemented in QChem (version 6.0.2) [34].
@@
We conducted LR-TD-DFT calculations using the Tamm–Dancoff approximation (TDA). Although TDA may not predict oscillator strengths as accurately as the random-phase approximation (RPA), RPA predictions of triplet energies are sensitive to triplet stability [35]. Therefore, TDA is preferred for automated high-throughput calculations, and previous work has shown that its STG predictions for azaphenalines are more accurate than that of RPA[21]. The performance of various DFAs for predicting STGs in azaphenalines has been discussed in prior studies [36, 37, 38].
@@
This study explores several DFAs spanning all levels of the Jacob’s ladder [39], namely, local density approximation (LDA): VWN5 [40] and PWLDA generalized gradient approximation (GGA): BP86, BLYP, PW91 [41, 42], PBE [43], mPWPW, mPWLYP hybrid GGA: B3LYP [44], PBE0 [45], X3LYP [46], mPW1PW, mPW1LYP, BHandHLYP [47] meta-GGA: TPSS [48], TPSSh [48], TPSS0 [48], M06L [49] hybrid meta-GGA: M062X [50] range-separated hybrids: $\omega$B97, [51] $\omega$B97X, [51]$\omega$B97X-D3, CAM-B3LYP [52], LC-BLYP [53], LC-PBE [54] double hybrid: $\omega$B97X-2 [55], B2PLYP [56], mPW2PLYP [57], B2GP-PLYP [58], PWPB95 [23], PBE-QIDH [59], PBE0-DH [60], $\omega$B2PLYP [61], $\omega$B2GP-PLYP [61], $\omega$B88PP86 [62], $\omega$PBEPP86 [62], SCS/SOS-B2PLYP21 [62], SCS-PBE-QIDH [62], SCS-B2GP-PLYP21 [62], SOS-B2GP-PLYP21 [62], SCS-$\omega$B2GP-PLYP21 [62], SOS-$\omega$B2GP-PLYP21 [62], SCS-RSX-QIDH [62], SOS-RSX-QIDH [62], SCS-$\omega$B88PP86 [62], SOS-$\omega$B88PP86 [62], SCS-$\omega$PBEPP86 [62], SOS-$\omega$PBEPP86 [62], DSD-BLYP [63], DSD-PBEP86 [63], DSD-PBEB95 [63], RSX-QIDH [64], and RSX-0H [65]. All DFT calculations were performed using Orca (version 6.0.0) [66, 67].
@@ 3 Bias Correction for Low-Variance Predictors
    The error metrics are defined as follows. Mean signed error (MSE): ${\rm MSE}=\mathbb{E}[\hat{y}_{1}-\hat{y}_{\rm ref}]$, mean absolute deviation (MAD): ${\rm MAD}=\mathbb{E}[|\hat{y}_{1}-\hat{y}_{\rm ref}|]$, standard deviation of error (SDE) for a sample of size $N$: ${\rm SDE}(\hat{y}_{1})=\sqrt{\frac{1}{N-1}\sum_{k=1}^{N}(\hat{y}_{1,k}-\hat{y}_{{\rm ref},k})^{2}}$. In the limit of a large sample (i.e., for a population), the square of SDE corresponds to the variance.
@@
    The term ‘bias’ refers to a systematic shift in a distribution and requires a reference for comparison. Throughout this study, bias specifically refers to systematic shifts in prediction error: $\hat{y}_{1}-\hat{y}_{\rm ref}$. Similarly, we distinguish between: ${\rm Var}(\hat{y}_{1})$ and ${\rm Var}(\hat{y}_{1}-\hat{y}_{\rm ref})$, where the former indicates spread around the mean while latter quantifies the spread of prediction errors, which is the primary focus of this study.
@@
This implies that the variance of $\hat{y}_{1}^{*}-\hat{y}^{\rm ref}$ is not simply $a^{2}$ times the variance of $\hat{y}_{1}-\hat{y}^{\rm ref}$, but is influenced by the distribution of the reference itself. Thus, only when $a=1$, the variance of $\hat{y}_{1}^{*}-\hat{y}^{\rm ref}$ exactly matches that of $\hat{y}_{1}-\hat{y}^{\rm ref}$. If $a=-1$, while the spread of $\hat{y}_{1}^{*}$ remains unchanged from the spread of $\hat{y}_{1}$, the spread of errors will increase. For general values of $a$, it is possible that fortuitously, the variance of prediction errors decreases after scaling. .
@@
In the above discussion, we designated the predictor with low MSE as an internal reference predictor, $\hat{y}_{2}$, interpreting it as the lower-bias model. However, this predictor will also have a low MAD only if all errors, $\hat{y}_{2}-\hat{y}^{\rm ref}$, have the same sign. In particular, if all errors are negative, then $\text{MAD}=-\text{MSE}$. In practice, a vanishing MSE for an approximate model often suggests a multimodal error distribution centered around zero, where positive and negative errors cancel each other out. In such cases, MSE can significantly underestimate the actual bias, making it an unreliable metric. A more robust measure of bias is MAD, which accounts for the magnitude of errors irrespective of sign. Thus, in this study, we select $\hat{y}_{2}$ as the predictor with the smallest MAD instead of the smallest MSE. This change does not affect any of the derivations presented above. Notably, the MSE of the scaled low-variance predictor, $\hat{y}_{1}^{*}$, remains close to that of $\hat{y}_{2}$, ensuring bias correction. Furthermore, since the variance of $\hat{y}_{1}^{*}$ remains approximately the same as that of $\hat{y}_{1}$ (for $a\approx 1$) the spread of errors is expected to be narrower. As a result, the MAD of $\hat{y}_{1}^{*}$ will typically be smaller than that of $\hat{y}_{2}$, reflecting a better balance between bias and variance.
@@ 4.1 Overall trends across methods
Figure 2 presents an overview of the performance of various theoretical methods using the TBE values from [14] as a reference. For consistency, all results are reported with the aug-cc-pVDZ basis set; results for other basis sets are provided in the Supplementary Information (SI), and the CC3 results are taken from [14]. Detailed error metrics for the wavefunction methods are provided in Table 1. For all methods listed in Table 1, increasing the basis set to aug-cc-pVTZ yields only minimal improvement, with EOM-CCSD being the sole exception, as it shows a slight deterioration in predictions with larger basis sets. The reference TBE S1 energies were computed at the CC3/aug-cc-pVTZ level with a post-CC3 correction via EOM-CCSDT (using the 6-31+G(d) basis set), while the T1 energies were obtained at the EOM-CCSD/aug-cc-pVTZ level with a post-CCSD correction via CC3 using the aug-cc-pVDZ basis set [14]. Although the TBE approach treats S1 and T1 energies using different protocols, both are estimated at higher levels of theory currently available.
Among all methods, CC3/aug-cc-pVDZ shows the lowest values for MSE, MAD, and SDE ($-1$, 10, and 13 meV (milli-eV), respectively; see Table 1). When combined with the aug-cc-pVTZ basis set, CC3 essentially reproduces the TBE results. The next best performance is observed for L-CC2 with an MAD and SDE of 16 and 13 meV. Overall, when using our more accurate value for system-10, ADC(2)/aug-cc-pVTZ yields an MAD and SDE of 20 and 16 meV, compared to 30 and 35 meV from Ref. [14]. Our L-ADC(2)/aug-cc-pVTZ results show an MAD and SDE of 22 and 13 meV, respectively, which are similar to the corresponding ADC(2) values of 20 and 16 meV.
Both the mean error and the spread of the error are significant for EOM-CCSD (Figure 2). Its MSE indicates that its predictions can serve as an upper bound for negative STGs; however, an SDE of approximately 80–90 meV diminishes confidence in predictions of gaps smaller than 0.1 eV, potentially leading to false negatives (i.e., systems with truly negative gaps might be predicted as having positive gaps). The large SDE suggests that the error is not systematic, so not all small, positive STGs predicted by EOM-CCSD can be reliably classified as negative. In particular, the similarity-transformed local variant, DLPNO-STEOM-CCSD, exhibits a bias exceeding 0.25 eV with an SDE around 0.1 eV (see Figure 2), rendering it unsuitable for applications involving negative STG systems [68, 69].
@@
Among the DFT methods, the dh-DFT approaches B2GP-PLYP and PBE-QIDH display very small SDEs, indicating that their deviations from the TBE values are primarily systematic. The spin-component-scaled (SCS) and opposite-spin-scaled (SOS) variants of these methods exhibit reduced MSE and MAD at the expense of larger SDEs. With few exceptions, the remaining DFAs show large errors, rendering them unreliable for first-principles identification of negative STG systems.

### v2

@@ 1 Introduction
The energy criterion for organic light-emitting diodes (OLEDs) to operate via thermally activated delayed fluorescence (TADF) is that the energy of the first excited singlet state (S1) must be no more than $0.1$ eV higher than that of the triplet state (T1)[1, 2]. This condition facilitates thermally-aided reverse intersystem crossing (RISC), enabling the theoretical complete transfer of the T1 population to the emissive state, S1[3]. A derivative of the triangular molecule, heptazine (aka heptaazaphenalene, 7AP), with a central nitrogen (N) atom in an anti-aromatic 12-annulene framework, showed delayed fluorescence without thermal assistance, indicating a negative S1–T1 gap (STG)[4]. This mechanism paves the way for designing OLEDs that leverage exothermic delayed fluorescence from inverted singlet and triplet excited states (DFIST). Direct spectroscopic evidence of a negative STG ($-0.047\pm 0.007$ eV) was obtained for pentaazaphenalene (5AP) using anion photoelectron spectroscopy and fluorescence measurements [5]. Negative STG was also was also observed in the transient photoluminescence spectrum of dialkylamine-substituted 5AP[6]. Computational studies have further suggested that azaphenalenes (APs) with other substitution patterns, as well as the boron (B) analog of azaphenalene (1AP), boraphenalene (1BP), also exhibit negative STGs[7, 8, 9, 10, 11, 12, 13, 14]. Non-alternant hydrocarbons and their substituted analogs show the potential for negative STGs[15, 16, 17, 18, 19]. Notably, substituted analogs of the non-fused bicyclic hydrocarbon have demonstrated negative STGs, attributed to through-bond charge-transfer states[20].
@@
A comprehensive search across $12,880$ structurally diverse small molecules revealed no exceptions to Hund’s rule, indicating that achieving the electronic structure criteria for ${\rm STG}<0$ requires nontrivial molecular structures[21]. Moreover, the typical magnitudes of negative STGs are much smaller or similar to the average errors associated with popular quantum chemistry and density functional theory approximations (DFAs). Consequently, due to the rarity of molecules with negative STGs and the limitations of current quantum chemical approximations, accurately identifying molecules exhibiting ${\rm STG}<0$ via high-throughput screening—with minimal false-positive and false-negative predictions—remains challenging. The linear-response time-dependent density functional theory (LR-TD-DFT) formalism provides qualitative accuracy—i.e., gets the sign of the STG for DFIST candidates correct—only within the double-hybrid (dh) DFT framework containing many-body correlation via the second-order perturbation theory (MP2). With the right combination of ingredients, dh-DFT can reliably predict a wide range of molecular properties[22, 23, 24, 23]. Nevertheless, selecting suitable exchange-correlation (XC) functionals within the dh-DFT framework remains challenging, especially for novel problems lacking sufficient reference data.
@@
We present a scheme that leverages the bias-variance tradeoff in the dh-DFT formalism. We evaluate various computational methods for predicting twelve STGs in triangular molecules, benchmarking them against previously reported theoretical best estimates[14]. We examine dh-DFAs exhibiting very low variance for STG predictions, and reparametrize them to minimize the bias. These optimal low-bias models are used as internal references to correct the predictions of the low-variance models. This internally referenced scaling approach is demonstrated by predicting STGs with low bias and low variance, resulting in excellent overall accuracy. We discuss the merits and drawbacks of this approach for its further applications.
@@ 2 Dataset and Computational Methods
We consider twelve systems (see Figure 1) using equilibrium geometries from Ref. [14], determined at the CCSD(T) level with the frozen core approximation and the cc-pVTZ basis set. Using these geometries, we performed single-point excited state calculations with various methods. In addition, we collected TBE of the S1 and T1 energies from Ref. [14] to serve as reference values for benchmarking. As stated in Ref. [14], structure 11 in Figure 1 (with $D_{3\rm h}$ symmetry) corresponds to a saddle point on the potential energy surface at the MP2/6-311G($d$,$p$) level, whereas the minimum energy geometry exhibits $C_{3\rm h}$ symmetry (structure 12). Similarly, a separate study[25] found that the high-symmetry forms of structures 6, 7, and 8 are not true minima and structure 2 is a very shallow potential well at the CCSD(T)/cc-pVTZ level[25]. Overall, four of the twelve structures are transition states; however, we include them in our benchmarking of STGs due to the availability of TBE-level results and the lack of alternative high-quality data.
@@
Laplace-transformed, density-fitted, local versions of correlated methods—CC2 [26] and ADC(2) [27]—offer attractive speedups for modeling excited states of medium-sized molecules. We denote these methods as L-CC2 and L-ADC(2), and performed calculations with the cc-pVDZ, aug-cc-pVDZ, cc-pVTZ, and aug-cc-pVTZ basis sets, using the corresponding JKFIT and MP2FIT auxiliary basis sets. L-CC2 and L-ADC(2) are the same methods denoted as LT-DF-CC2 and LT-DF-ADC(2) elsewhere. L-CC2 and L-ADC(2) calculations were conducted using Molpro (version 2015.1) [28]. L-CC2 has been shown to deviate from the canonical variant CC2 by $\leq 0.05$ eV, irrespective of the character of the excited states, including that of charge-transfer states for which CC2 can have huge error [29]. The systems studied in this work do not have charge-transfer characteristics, hence CC2 delivers good accuracy for STGs in agreement with TBEs[14]. L-CC2 has demonstrated excellent accuracy in modeling the excited states of BODIPY derivatives [30, 31]. For comparison, we collected S1 and T1 energies obtained using the linear-response coupled-cluster methods, EOM-CCSD and CC3, with the aug-cc-pVDZ basis set from Ref. [14]. Additionally, we performed ADC(2) calculations using the cc-pVDZ, cc-pVTZ, aug-cc-pVDZ, and aug-cc-pVTZ basis sets with the resolution-of-the-identity (RI) approximation [32, 33] as implemented in QChem (version 6.0.2) [34].
@@
We conducted LR-TD-DFT calculations using the Tamm–Dancoff approximation (TDA). Although TDA may not predict oscillator strengths as accurately as the random-phase approximation (RPA), RPA predictions of triplet energies are sensitive to triplet stability [35]. Therefore, TDA is preferred for automated high-throughput calculations, and previous work has shown that its STG predictions for azaphenalenes are more accurate than that of RPA[21]. The performance of various DFAs for predicting STGs in azaphenalines has been discussed in prior studies [36, 37, 38].
@@
This study explores several DFAs spanning all levels of the Jacob’s ladder [39], namely, local density approximation (LDA): VWN5 [40] and PWLDA; generalized gradient approximation (GGA): BP86, BLYP, PW91 [41, 42], PBE [43], mPWPW, and mPWLYP; hybrid GGA: B3LYP [44], PBE0 [45], X3LYP [46], mPW1PW, mPW1LYP, and BHandHLYP [47]; meta-GGA: TPSS [48], TPSSh [48], TPSS0 [48], and M06L [49] hybrid meta-GGA: M062X [50]; range-separated hybrids: $\omega$B97, [51] $\omega$B97X, [51]$\omega$B97X-D3, CAM-B3LYP [52], LC-BLYP [53], and LC-PBE [54]; double hybrid: $\omega$B97X-2 [55], B2PLYP [56], mPW2PLYP [57], B2GP-PLYP [58], PWPB95 [23], PBE-QIDH [59], PBE0-DH [60], $\omega$B2PLYP [61], $\omega$B2GP-PLYP [61], $\omega$B88PP86 [62], $\omega$PBEPP86 [62], SCS/SOS-B2PLYP21 [62], SCS-PBE-QIDH [62], SCS-B2GP-PLYP21 [62], SOS-B2GP-PLYP21 [62], SCS-$\omega$B2GP-PLYP21 [62], SOS-$\omega$B2GP-PLYP21 [62], SCS-RSX-QIDH [62], SOS-RSX-QIDH [62], SCS-$\omega$B88PP86 [62], SOS-$\omega$B88PP86 [62], SCS-$\omega$PBEPP86 [62], SOS-$\omega$PBEPP86 [62], DSD-BLYP [63], DSD-PBEP86 [63], DSD-PBEB95 [63], RSX-QIDH [64], and RSX-0H [65]. All DFT calculations were performed using Orca (version 6.0.0) [66, 67].
@@ 3 Bias Correction for Low-Variance Predictors
    The error metrics are defined as follows. Mean signed error (MSE): ${\rm MSE}(\hat{y}_{1})=\mathbb{E}[\hat{y}_{1}-\hat{y}_{\rm ref}]$, mean absolute deviation (MAD): ${\rm MAD}(\hat{y}_{1})=\mathbb{E}[|\hat{y}_{1}-\hat{y}_{\rm ref}|]$, standard deviation of error (SDE) for a sample of size $N$: ${\rm SDE}(\hat{y}_{1})=\sqrt{\frac{1}{N-1}\sum_{k=1}^{N}(\hat{y}_{1,k}-\hat{y}_{{\rm ref},k})^{2}}$. In the limit of a large sample (i.e., for a population), the square of SDE corresponds to the variance.
@@
    The term ‘bias’ refers to a systematic shift in a distribution and requires a reference for comparison. Throughout this study, bias specifically refers to systematic shifts in prediction error: $\hat{y}_{1}-\hat{y}_{\rm ref}$. Similarly, we distinguish between: ${\rm Var}(\hat{y}_{1})$ and ${\rm Var}(\hat{y}_{1}-\hat{y}_{\rm ref})$, where the former indicates spread around the mean value of $\hat{y}_{1}$ while latter quantifies the spread of its prediction errors, which is the primary focus of this study.
@@
This implies that the variance of $\hat{y}_{1}^{*}-\hat{y}^{\rm ref}$ is not simply $a^{2}$ times the variance of $\hat{y}_{1}-\hat{y}^{\rm ref}$, but is influenced by the distribution of the reference itself. Thus, only when $a=1$, the variance of $\hat{y}_{1}^{*}-\hat{y}^{\rm ref}$ exactly matches that of $\hat{y}_{1}-\hat{y}^{\rm ref}$. If $a=-1$, while the spread of $\hat{y}_{1}^{*}$ remains unchanged from that of $\hat{y}_{1}$, the spread of the errors of $\hat{y}_{1}^{*}$ will increase. For general values of $a$, it is possible that fortuitously, the variance of prediction errors decreases after scaling.
@@
In the above discussion, we designated the predictor with low MSE as an internal reference, $\hat{y}_{2}$, interpreting it as the lower-bias model. However, this predictor will also have a low MAD only if all errors, $\hat{y}_{2}-\hat{y}^{\rm ref}$, have the same sign. In particular, if all errors are negative, then $\text{MAD}=-\text{MSE}$. In practice, a vanishing MSE for an approximate model often suggests a multimodal error distribution centered around zero, where positive and negative errors cancel each other out. In such cases, MSE can significantly underestimate the actual bias, making it an unreliable metric. A more robust measure of bias is MAD, which accounts for the magnitude of errors irrespective of sign. Thus, in this study, we select $\hat{y}_{2}$ as the predictor with the smallest MAD instead of the smallest MSE. This change does not affect any of the derivations presented above. Notably, the MSE of the scaled low-variance predictor, $\hat{y}_{1}^{*}$, remains close to that of $\hat{y}_{2}$, ensuring bias correction. Furthermore, since the variance of $\hat{y}_{1}^{*}$ remains approximately the same as that of $\hat{y}_{1}$ (for $a\approx 1$) the spread of errors is expected to be narrower. As a result, the MAD of $\hat{y}_{1}^{*}$ will typically be smaller than that of $\hat{y}_{2}$, reflecting a better balance between bias and variance.
@@ 4.1 Overall trends across methods
Figure 2 presents an overview of the performance of various theoretical methods using the TBE of STGs from Ref. [14] as a reference. For consistency, all results are reported with the aug-cc-pVDZ basis set; results for other basis sets are provided in the Supplementary Information (SI). CC3 results are taken from Ref. [14]. Detailed error metrics for the wavefunction methods are provided in Table 1. For all methods listed in Table 1, increasing the basis set to aug-cc-pVTZ yields only minimal improvement. The reference (i.e., TBE) S1 energies were computed at the CC3/aug-cc-pVTZ level with a post-CC3 correction via EOM-CCSDT (using the 6-31+G(d) basis set), while the T1 energies were obtained at the EOM-CCSD/aug-cc-pVTZ level with a post-CCSD correction via CC3 using the aug-cc-pVDZ basis set [14]. Although the TBE approach treats S1 and T1 energies using different protocols, both are estimated at higher levels of theory currently available.
Among all methods, CC3/aug-cc-pVDZ shows the lowest values for MSE, MAD, and SDE ($-1$, 10, and 13 meV (milli-eV), respectively; see Table 1). When combined with the aug-cc-pVTZ basis set, CC3 essentially reproduces the TBE results. The next best performance is observed for L-CC2 with an MAD and SDE of 16 and 13 meV. L-ADC(2)/aug-cc-pVTZ results show an MAD and SDE of 22 and 13 meV, respectively, which are similar to the corresponding ADC(2) values of 20 and 16 meV.
Both the mean error and the spread of the error are significant for EOM-CCSD (Figure 2). Its MSE indicates that its predictions can serve as an upper bound for negative STGs; however, an SDE of approximately 80–90 meV diminishes confidence in predictions of gaps smaller than 0.1 eV, potentially leading to false negatives (i.e., systems with truly negative gaps might be predicted as having positive gaps). The large SDE suggests that the error is not systematic, hence, not all small, positive STGs predicted by EOM-CCSD may be reliably expected to be negative. In particular, the similarity-transformed local variant, DLPNO-STEOM-CCSD, exhibits a bias exceeding 0.25 eV with an SDE around 0.1 eV (see Figure 2), rendering it unsuitable for applications involving negative STG systems [68, 69].
@@
Among the DFT methods, the dh-DFT approaches B2GP-PLYP and PBE-QIDH display very small SDEs (Figure 2), indicating that their deviations from the TBE values are primarily systematic. The spin-component-scaled (SCS) and opposite-spin-scaled (SOS) variants of these methods exhibit reduced MSE and MAD at the expense of larger SDEs. With few exceptions, the remaining DFAs show large errors, rendering them unreliable for first-principles identification of negative STG systems.